Prove that any polynomial of odd degree has atleast one real root


1
Expert's answer
2021-07-21T11:52:32-0400

Solution:

We want to show that if "P(x)=a_{n} x^{n}+a_{n-1} x^{n-1}+\cdots+a_{1} x+a_{0}" is a polynomial with n odd and "a_{\mathrm{n}} \neq 0" , then there is a real number c, such that P(c)=0

Firstly we know that every polynomial is continuous on the real line. We also know that

"\lim _{n \rightarrow \infty} \frac{P(x)}{a_{n} x^{n}}=1 \text { and } \lim _{n \rightarrow-\infty} \frac{P(x)}{a_{n} x^{n}}=1"

Consequently for "|x|" large enough, "P(x)" and "a_nx^n" have the same sign. But "a_nx^n" has opposite signs for positive "x" and negative "x". Thus it follows that if "a_n>0", there are real numbers "x_0<x_1" such that "P(x_0)<0" and "P(x_1)>0". Similarly, if "a_n<0", we can find "x_0<x_1" such that "P(x_0)>0" and "P(x_1)<0". In each case, it now follows directly from the Intermediate Value Theorem that (for d = 0) there is a real number "c\in\left[x_{0}, x_{1}\right]\ with\ P(c)=0"